OpenStudy (anonymous):

a student finishes and exam in 2/3 the times it takes him to finish the second half. If the entire exam takes him an hour, how many monutes does he spend on the first exam a. 20 b. 24 c. 27 d 36

OpenStudy (anonymous):

I presume he spent 2/3 of the time on the second half? If so, just take the number of minutes, 60, and multiply it by 2/3 to get 40 minutes. That's how long the student spent on the second half, now subtract 40 from 60 to get your answer.

OpenStudy (anonymous):

Let t be the completion time for the second half. Solve the following for t:\[\frac{2}{3} t+t= 60\]t = 36 (2/3)*36 = 24 -> (b.)

OpenStudy (anonymous):

You need to write these equations: 1) total time = firstHalf + secondHalf 2) firstHalf = (2/3) * secondHalf 3) total time = 60 minutes Then it is a matter of solving this system of equations using substitution. The attached file shows the steps.

OpenStudy (anonymous):

20 min

OpenStudy (anonymous):

thx touchy math and moriaw answers was well understood